s Consider the differentiad eqation (a) Find the characteristic equation, determine its roots and their corresponding multipl can anyone answer this
Answer #1

yt y=0 characteristic equation a) is m (m3-1) = n (m-i) (m? 7 m +1)=0 [23-63 = Ca-b) (ahtab +67 m=o (m-1)=0 = m = 1 m +m m=c) of m, ma = &#ip, m3, ma then = edx [cicospx + asinpx] + e + cet Mzo na Therefore, the general solution is Xx) = é 12x [cic​​​​​​